OpenStudy (anonymous):

The circumference of a circle is 60 pi. what is the length of an arc of 140 degrees? help please

OpenStudy (anonymous):

ok .... 60 =2pi r r = 30/pi s = r *theta s = 140 * 30 /pi= 1336.90

OpenStudy (anonymous):

r should be 30. 60pi/2pi

OpenStudy (anonymous):

oh thnx i forgot pi

OpenStudy (anonymous):

so your answer would be s = 140 *30

OpenStudy (anonymous):

Convert degrees to radians. 140 degrees x pi/180 L = (140)(pi)(30) / 180 = 73.27